Please help me with this:

I've asked this same question the first time, but no tutor helped me yet!

A pole 24ft high has a shadow 8ft long. A nearby pole is 72ft tall. How long is the shadow?

I need an immediate answer please?

Answer by jim_thompson5910(35256)

24/8 = 72/x

24*x = 8*72 ... cross multiply

24x = 576

x = 576/24

x = 24

The shadow is 24 feet long.